### v1

@@ 1 Introduction
An important task for string theory is to make connections to observational physics. One of the most promising places where such links may be made is in the context of cosmology. The normal approach to this starts through dimensional reduction of string theory into a 4-dimensional effective field theory; all actual calculations are then performed within this 4-dimensional theory and it is only rarely that the underlying 10-dimensional theory is considered explicitly. In this respect, a detailed recent review of string cosmology is Cicoli:2023opf (see Baumann:2009ni ; Brandenberger:2023ver for other reviews with a slightly different focus).
One potential non-standard cosmological epoch in the early universe is that of a kination epoch (the name first appears in Joyce:1996cp ), an epoch dominated by the kinetic energy of a rolling scalar field. Such epochs naturally arise in string theory when moduli fields roll down steep potentials, for example when running from the centre of moduli space to the boundaries (a recent detailed analysis of this is Apers:2024ffe ; see Conlon:2022pnx ; Revello:2023hro ; Shiu:2023fhb ; Shiu:2023nph ; Seo:2024qzf for some other recent studies of kination epochs in string theory).
One intriguing aspect of kination epochs in the context of IIB string compactifications, realised in Apers:2022cyl , is that the epoch also admits a direct 10-dimensional description in terms of a higher-dimensional Kasner solution. This provides an interesting contrast to most work in string cosmology, where no such 10-dimensional description exists. This provides one of the few examples where it is possible to go beyond the 4d description and work directly within the 10d theory (another famous example where a 10d description exists is the GKP stage of IIB flux compactifications Giddings:2001yu , prior to the step of Kähler moduli stabilisation for which Kachru:2003aw and Balasubramanian:2005zx are the best-known examples).
The purpose of this note is to build on the analysis of Apers:2022cyl and explore this link further. There are three sections. In the first, we extend the results of Apers:2022cyl to other string theories beyond the IIB orientifold examples considered there. In the second, we fill in more details on the perturbations of the 10d theory and their connection to moduli backgrounds in the 4d kination theory, going beyond the overall Kasner dynamics associated to the volume mode that was considered in Apers:2022cyl . In the final one, we explore the consequence of a gas of winding modes during the kination epoch.
@@ 2 Kination In General String Theories with Multiple Rolling Moduli
We first consider kination epochs in more general string scenarios than the type IIB orientifold examples with just a single rolling volume modulus considered in Apers:2022cyl . As a single framework to capture multiple examples, we consider a 4d EFT which contains two moduli, a volume modulus and a dilaton, both with vanishing potential. [^1] Even if the potential does not strictly vanish, if it is steep enough, fields enter a kination epoch as they roll down it and the potential can effectively be neglected. The kinetic energy for these two moduli are determined by a Kähler potential which takes the form
@@
With an appropriate redefinition of the time coordinate, we will confirm that this is equivalent to a Mueller metric MUELLER199037 which is a generalisation of the Kasner metric of the form
@@ 2.1 Kination and its 10d uplift in type IIA string theory
Here we make this connection explicit for type IIA string theory, where the dilaton and volume moduli $s$ and $u$ are defined by (for example, see Grimm:2004ua )
@@ 3 Perturbations of 10d Kasner Solutions and their relation to 4d Moduli Backgrounds
Given the interest of Kasner solutions as a 10d completion of kination solutions, it is worthwhile understanding their perturbations better (the paper Sugiyama:2020mbu describes gravitational waves in the context of Kasner spacetimes including the Unruh effect). From a 4-d perspective, such perturbations correspond to the presence of additional matter/radiation content in the 4d theory.
@@
We start with an analysis of the 10d solution. Anticipating section 3.2 where we relate perturbations of the Kasner metric to moduli backgrounds in the kination solution, it is useful to express 10d quantities in terms of the kination conformal time $\eta$ [^2] As field perturbations in 4d theories are most naturally expressed in terms of the conformal time. This relates to the time $t$ of the 10-dimensional Kasner solution through $\eta=\tfrac{3}{4}t^{4/3}$, in which case the Kasner metric takes the form
@@
where $H_{ij}$ is now an amplitude matrix. Equation 29 requires that $\operatorname{tr}H=-(\Sigma_{\alpha}m_{\alpha})A$, while Equation 30 implies that $H_{ij}k_{j}=0$. This shows that perturbations in the volume modulus are necessarily accompanied by perturbations in the noncompact part of the metric. During compactification and dimensional reduction, these noncompact perturbations correspond to the 4d dilaton, which is then re-absorbed into the final definition of the noncompact Einstein frame metric.
@@
$\nabla^{2}f_{i}-\partial_{\eta}^{2}f_{i}-\sum_{j}\partial_{i}\partial_{j}f_{j}$ $=$ $0$ (36)
@@
As a check on the above three types of perturbation, we can count the number of degrees of freedom. In the temporal gauge there are a total of 45 free metric components. The equations of motion impose six polarization conditions on the radiation perturbations $\boldsymbol{A}_{\alpha}\cdot\boldsymbol{k}=0$, a single trace condition on volume perturbations $\sum_{\alpha}m_{\alpha}=-\operatorname{tr}H$, and three polarization conditions on the noncompact gravitational waves $k_{i}H_{ij}=0$. This leaves $45-6-3-1=35$ degrees of freedom, matching the number of degrees of freedom of a graviton in ten dimensions.
@@
When expanding around a 10-dimensional background metric $\bar{g}_{MN}$ that solves the Einstein equation as $g_{MN}=\bar{g}_{MN}+H_{MN}$, then up to cubic contributions $\mathcal{O}(H^{3})$ the Einstein–Hilbert action can be rewritten as the Fierz–Pauli action in curved spacetime Fierz:1939 ; Ortin:2015
$\begin{split}S=-\frac{1}{2\kappa^{2}}\int d^{10}x\sqrt{-\bar{g}}\bigg{[}&\frac{1}{4}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{M}H^{AB}-\frac{1}{2}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{A}H^{MB}\\ &+\frac{1}{2}\bar{\nabla}_{M}H^{MN}\bar{\nabla}_{N}H-\frac{1}{4}\bar{\nabla}_{M}H\bar{\nabla}^{M}H\bigg{]},\end{split}$ (40)
where $H=H_{A}{}^{A}$. In this case, the background metric is the Kasner metric. The bar over the covariant derivative indicates that this is the ten dimensional covariant derivative as opposed to the four dimensional covariant derivative that will appear after compactification.
@@
where $\phi=\sum_{i}\phi_{ii}$, $g_{\mu\nu}=\eta_{\mu\nu}\zeta^{-1}$, and $\zeta=\sqrt{4\eta/3}$. The Kasner compactification differs from a standard compactification on a torus by the time-dependent factors of $\zeta$ Han:1998sg .
@@
$\frac{\mathcal{L}}{\sqrt{-g}}$ $\supset$ $\frac{2}{9}\zeta^{2}A^{2}-\frac{2}{3}\zeta^{4}A^{\nu i}\nabla_{0}A_{\nu i}$ (42)
$=$ $-\frac{2}{9}\zeta^{2}A^{2}-\frac{2}{3}\bigg{(}\frac{1}{2}\nabla_{0}(\zeta^{4}A^{\nu i}A_{\nu i})-\zeta^{2}A^{2}\bigg{)}$ (43)
$=$ $-\frac{2}{9}\zeta^{2}A^{2}-\frac{2}{3}\bigg{(}\frac{1}{2}\nabla_{\mu}X^{\mu}-\frac{1}{3}\zeta^{2}A^{2}\bigg{)}$ (44)
$=$ $-\frac{1}{3}\nabla_{\mu}X^{\mu},$ (45)
@@
$\begin{split}\mathcal{L}=\sqrt{-g}\mathcal{V}_{6,0}\zeta^{3}\bigg{[}&\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h-\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h+\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}-\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}\\ &-\tfrac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\tfrac{1}{4}\zeta^{2}g^{ij}F_{i,\mu\nu}F^{\mu\nu}_{j}-\zeta^{2}g^{ik}g^{jp}\nabla_{\mu}\phi_{ij}\nabla^{\mu}\phi_{kp}\\ &-\tfrac{4}{3}\zeta^{-1}\phi\nabla_{0}h+\zeta^{-1}h_{0}^{\mu}\nabla_{\mu}h-\tfrac{8}{3}\zeta^{-1}h_{0}^{\mu}\nabla_{\mu}\phi\bigg{]},\end{split}$ (46)
where $h=h^{\mu}{}_{\mu}$ and $F_{i,\mu\nu}$ are the field-strength tensors for the $A_{i,\mu}$ gauge fields. To go into the Einstein frame requires performing a Weyl transformation on the metric $g_{\mu\nu}\rightarrow\mathcal{V}_{6}^{-1}g_{\mu\nu}$ and its perturbations $h_{\mu\nu}$, giving
$\begin{split}S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}&\bigg{[}\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h-\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h+\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}-\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}\\ &-\tfrac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\tfrac{1}{4\tilde{g}^{2}}\sum_{i=1}^{6}F_{i}^{2}-\sum_{ij}\partial_{\mu}\phi_{ij}\partial^{\mu}\phi_{ij}\\ &-\tfrac{4}{3}\phi B^{\mu}\partial_{\mu}h+B^{\mu}h_{\mu}^{\nu}\partial_{\nu}h-\tfrac{8}{3}B^{\mu}h^{\nu}_{\mu}\partial_{\nu}\phi\bigg{]},\end{split}$ (47)
where $B^{\mu}=(\mathcal{V}_{6,0}^{-1}\zeta^{-4},0,0,0)$ is a non-dynamical background vector field that goes to zero at late times and $\tilde{g}^{2}=\zeta^{-1}\mathcal{V}^{-1}_{6}\sim\eta^{-2}$ is a time-dependent coupling for the gauge field.
@@
$\begin{split}S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}&\bigg{[}\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h-\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h+\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}-\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}\\ &-\tfrac{1}{2}\partial_{\mu}\varphi\partial^{\mu}\varphi-\tfrac{1}{4\tilde{g}^{2}}\sum_{i=1}^{6}F_{i}^{2}-\tfrac{1}{2}\sum_{i=1}^{5}\partial_{\mu}\Delta_{i}\partial^{\mu}\Delta_{i}-\sum_{i\neq j}\partial_{\mu}\phi_{ij}\partial^{\mu}\phi_{ij}\\ &-\tfrac{2}{3}\varphi B^{\mu}\partial_{\mu}h+B^{\mu}h_{\mu}^{\nu}\partial_{\nu}h-\tfrac{4}{3}B^{\mu}h^{\nu}_{\mu}\partial_{\nu}\varphi\bigg{]}.\end{split}$ (49)
At small scales or at late times the last line in action goes to zero since $B^{\mu}\rightarrow 0$, while the first line reduces to the leading order perturbation of the Ricci scalar around a flat metric since the covariant derivatives reduce to regular derivatives. This corresponds to letting the Hubble scale become small relative to the energy scale associated with the moduli fields in which case the perturbations behave like perturbations in a standard toroidal compactification.
@@
The ten dimensional gauge transformation of the Fierz–Pauli action
$H_{MN}\rightarrow H_{MN}+\nabla_{M}\xi_{N}+\nabla_{N}\xi_{M},$ (50)
results, after the Weyl transformation, in the simultaneous gauge transformations
$\begin{split}h_{\mu\nu}&\rightarrow h_{\mu\nu}+\zeta^{3}\nabla_{\mu}\xi_{\nu}+\zeta^{3}\nabla_{\nu}\xi_{\mu}-2\zeta^{-1}g_{\mu\nu}\xi_{0},\\ A_{\mu i}&\rightarrow A_{\mu i}+\zeta^{-1}\partial_{\mu}\xi_{i}-\tfrac{2}{3}\zeta^{-3}\delta_{\mu 0}\xi_{i},\\ \varphi&\rightarrow\varphi-\tfrac{4}{\sqrt{3}}\zeta^{-1}\xi_{0}.\end{split}$ (51)
The $\xi_{i}$ components result in gauge transformations for the gauge field and leave the $F_{i}^{2}$ term invariant. Meanwhile, the $\xi_{\mu}$ transformations act nontrivially on the 4d dilaton and graviton. In the appendix, we illustrate the explicit gauge invariance of 4d action for one of these transformations.

### v2

@@ 1 Introduction
An important task for string theory is to make connections to observational physics. One of the most promising places where such links may be made is in the context of cosmology. The normal approach to this starts through dimensional reduction of string theory into a 4-dimensional effective field theory; all actual calculations are then performed within this 4-dimensional theory and it is only rarely that the underlying 10-dimensional theory is considered explicitly. In this respect, a detailed recent review of string cosmology is Cicoli et al. [2024] (see Baumann and McAllister [2009]; Brandenberger [2023] for other reviews with a slightly different focus).
One potential non-standard cosmological epoch in the early universe is that of a kination epoch (the name first appears in Joyce [1997]), an epoch dominated by the kinetic energy of a rolling scalar field. Such epochs naturally arise in string theory when moduli fields roll down steep potentials, for example when running from the centre of moduli space to the boundaries (a recent detailed analysis of this is Apers et al. [2024]; see Conlon and Revello [2022]; Revello [2024]; Shiu et al. [2023b, a]; Seo [2024] for some other recent studies of kination epochs in string theory).
One intriguing aspect of kination epochs in the context of IIB string compactifications, realised in Apers et al. [2023], is that the epoch also admits a direct 10-dimensional description in terms of a higher-dimensional Kasner solution. This provides an interesting contrast to most work in string cosmology, where no such 10-dimensional description exists. This provides one of the few examples where it is possible to go beyond the 4d description and work directly within the 10d theory (another famous example where a 10d description exists is the GKP stage of IIB flux compactifications Giddings et al. [2002], prior to the step of Kähler moduli stabilisation for which Kachru et al. [2003] and Balasubramanian et al. [2005] are the best-known examples).
The purpose of this note is to build on the analysis of Apers et al. [2023] and explore this link further. There are three sections. In the first, we extend the results of Apers et al. [2023] to other string theories beyond the IIB orientifold examples considered there. In the second, we fill in more details on the perturbations of the 10d theory and their connection to moduli backgrounds in the 4d kination theory, going beyond the overall Kasner dynamics associated to the volume mode that was considered in Apers et al. [2023]. In the final one, we explore the consequence of a gas of winding modes during the kination epoch.
@@ 2 Kination In General String Theories with Multiple Rolling Moduli
We first consider kination epochs in more general string scenarios than the type IIB orientifold examples with just a single rolling volume modulus considered in Apers et al. [2023]. As a single framework to capture multiple examples, we consider a 4d EFT which contains two moduli, a volume modulus and a dilaton, both with vanishing potential. [^1] Even if the potential does not strictly vanish, if it is steep enough, fields enter a kination epoch as they roll down it and the potential can effectively be neglected. The kinetic energy for these two moduli are determined by a Kähler potential which takes the form
@@
With an appropriate redefinition of the time coordinate, we will confirm that this is equivalent to a Mueller metric Mueller [1990] which is a generalisation of the Kasner metric of the form
@@ 2.1 Kination and its 10d uplift in type IIA string theory
Here we make this connection explicit for type IIA string theory, where the dilaton and volume moduli $s$ and $u$ are defined by (for example, see Grimm and Louis [2005])
@@ 3 Perturbations of 10d Kasner Solutions and their relation to 4d Moduli Backgrounds
Given the interest of Kasner solutions as a 10d completion of kination solutions, it is worthwhile understanding their perturbations better (the paper Sugiyama et al. [2021] describes gravitational waves in the context of Kasner spacetimes including the Unruh effect). From a 4-d perspective, such perturbations correspond to the presence of additional matter/radiation content in the 4d theory.
@@
We start with an analysis of the 10d solution. Anticipating section 3.2 where we relate perturbations of the Kasner metric to moduli backgrounds in the kination solution, it is useful to express 10d quantities in terms of the kination conformal time $\eta$. [^2] As field perturbations in 4d theories are most naturally expressed in terms of the conformal time. This relates to the time $t$ of the 10-dimensional Kasner solution through $\eta=\tfrac{3}{4}t^{4/3}$, in which case the Kasner metric takes the form
@@
where $H_{ij}$ is now an amplitude matrix. Eq. 29 requires that $\operatorname{tr}H=-(\Sigma_{\alpha}m_{\alpha})A$, while Eq. 30 implies that $H_{ij}k_{j}=0$. This shows that perturbations in the volume modulus are necessarily accompanied by perturbations in the noncompact part of the metric. During compactification and dimensional reduction, these noncompact perturbations correspond to the 4d dilaton, which is then re-absorbed into the final definition of the noncompact Einstein frame metric.
@@
$\nabla^{2}f_{i}-\partial_{\eta}^{2}f_{i}-\sum_{j}\partial_{i}\partial_{j}f_{j}$ $=$ $0.$ (36)
@@
As a check on the above three types of perturbation, we can count the number of degrees of freedom. In the temporal gauge there are a total of 45 free metric components. The equations of motion impose six polarization conditions on the radiation perturbations $\boldsymbol{A}_{\alpha}\cdot\boldsymbol{k}=0$, a single trace condition on volume perturbations $\sum_{\alpha}m_{\alpha}=-\operatorname{tr}H$, and three polarization conditions on the noncompact gravitational waves $k_{i}H_{ij}=0$. This leaves 45631=3545-6-3-1=35 degrees of freedom, matching the number of degrees of freedom of a graviton in ten dimensions.
@@
When expanding around a 10-dimensional background metric $\bar{g}_{MN}$ that solves the Einstein equation as $g_{MN}=\bar{g}_{MN}+H_{MN}$, then up to cubic contributions $\mathcal{O}(H^{3})$ the Einstein–Hilbert action can be rewritten as the Fierz–Pauli action in curved spacetime Fierz and Pauli,Wolfgang [1939]; Ortín [2015]
$\begin{split}S=-\frac{1}{2\kappa^{2}_{10}}\int d^{10}x\sqrt{-\bar{g}}\bigg[&\frac{1}{4}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{M}H^{AB}-\frac{1}{2}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{A}H^{MB}\\ &+\frac{1}{2}\bar{\nabla}_{M}H^{MN}\bar{\nabla}_{N}H-\frac{1}{4}\bar{\nabla}_{M}H\bar{\nabla}^{M}H\bigg],\end{split}$ (40)
where $H=H_{A}{}^{A}$. In this case, the background metric is the Kasner metric $g_{MN}=(g_{\mu\nu},g_{ij})=(\zeta^{-1}\eta_{\mu\nu},\zeta\delta_{ij})$, where $\zeta=\sqrt{4\eta/3}$. The bar over the covariant derivative indicates that this is the ten dimensional covariant derivative as opposed to the four dimensional covariant derivative that will appear after compactification.
@@
where $\phi=\delta^{ij}\phi_{ij}$. The Kasner compactification differs from a standard compactification on a torus by the time-dependent factors of $\zeta$ Han et al. [1999]. The expressions for the four terms in the Fierz–Pauli action in terms of the metric decomposition Eq. (41) are found in the appendix.
@@
$\frac{2\kappa_{10}^{2}}{\mathcal{V}_{6,0}\sqrt{-g}}\mathcal{L}$ $\supset$ $\frac{2}{9}\zeta^{2}A^{2}+\frac{2}{3}\zeta^{4}A^{\nu i}\nabla_{0}A_{\nu i}$ (42)
$=$ $\frac{2}{9}\zeta^{2}A^{2}+\frac{2}{3}\bigg(\frac{1}{2}\nabla_{0}(\zeta^{4}A^{\nu i}A_{\nu i})-\zeta^{2}A^{2}\bigg)$ (43)
$=$ $\frac{2}{9}\zeta^{2}A^{2}+\frac{2}{3}\bigg(\frac{1}{2}\nabla_{\mu}X^{\mu}-\frac{1}{3}\zeta^{2}A^{2}\bigg)$ (44)
$=$ $\frac{1}{3}\nabla_{\mu}X^{\mu},$ (45)
@@
$\begin{split}\mathcal{L}=\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\zeta^{3}\bigg[&\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h-\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h+\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}-\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}\\ &-\tfrac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\tfrac{1}{4}\zeta^{2}g^{ij}F_{i,\mu\nu}F^{\mu\nu}_{j}-\zeta^{2}g^{ik}g^{jp}\nabla_{\mu}\phi_{ij}\nabla^{\mu}\phi_{kp}\\ &-\tfrac{4}{3}\zeta^{-1}\phi\nabla_{0}h+\zeta^{-1}h_{0}^{\mu}\nabla_{\mu}h-\tfrac{8}{3}\zeta^{-1}h_{0}^{\mu}\nabla_{\mu}\phi\bigg],\end{split}$ (46)
where $h=h^{\mu}{}_{\mu}$ and $\kappa_{4}^{-2}=\mathcal{V}_{6,0}\kappa_{10}^{-2}$ is the 4d gravitational coupling constant. Meanwhile, $F_{i,\mu\nu}$ are the field-strength tensors for the abelian $A_{i,\mu}$ gauge fields. To go into the Einstein frame requires performing a Weyl transformation on the metric $g_{\mu\nu}\rightarrow\mathcal{\zeta}^{-3}g_{\mu\nu}$ and its perturbations $h_{\mu\nu}\rightarrow\zeta^{-3}h_{\mu\nu}$. The Weyl transformation rescales each term in the second and third lines individually to eliminate the $\zeta^{3}$ prefactor. The first line transforms to cancel this prefactor while also giving rise to two total derivatives that can be dropped and a term that cancels with the $h_{0}^{\mu}\nabla_{\mu}h$ term on the third line. The resulting action is
$\begin{split}S=\frac{1}{2\kappa^{2}_{4}}\int d^{4}x\sqrt{-g}&\bigg[\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h-\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h+\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}-\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}\\ &-\tfrac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\tfrac{1}{4\tilde{g}^{2}}\sum_{i=1}^{6}F_{i}^{2}-\sum_{ij}\partial_{\mu}\phi_{ij}\partial^{\mu}\phi_{ij}\\ &+\tfrac{1}{\sqrt{3}}\phi B^{\mu}\partial_{\mu}h+\tfrac{2}{\sqrt{3}}B^{\mu}h^{\nu}_{\mu}\partial_{\nu}\phi\bigg],\end{split}$ (47)
where we introduce the vector field $B^{\mu}=-\tfrac{4}{\sqrt{3}}(\zeta^{-4},0,0,0)$ and $\tilde{g}^{2}=\zeta^{-1}\mathcal{\zeta}^{-3}\sim\eta^{-2}$ is a time-dependent coupling for the gauge field.
@@
$\begin{split}S=\frac{1}{2\kappa^{2}_{4}}\int d^{4}x\sqrt{-g}&\bigg[\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h-\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h+\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}-\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}\\ &-\tfrac{1}{2}\partial_{\mu}\varphi\partial^{\mu}\varphi-\tfrac{1}{4\tilde{g}^{2}}\sum_{i=1}^{6}F_{i}^{2}-\tfrac{1}{2}\sum_{i=1}^{5}\partial_{\mu}\Delta_{i}\partial^{\mu}\Delta_{i}-\sum_{i\neq j}\partial_{\mu}\phi_{ij}\partial^{\mu}\phi_{ij}\\ &+\tfrac{1}{2}\varphi B^{\mu}\partial_{\mu}h+B^{\mu}h^{\nu}_{\mu}\partial_{\nu}\varphi\bigg].\end{split}$ (49)
The components of the action involving the metric perturbations $h_{\mu\nu}$ and the dilaton $\varphi$ match exactly the action found by expanding out the kination action with $\Phi=\bar{\Phi}+\varphi$ and $g_{\mu\nu}\rightarrow g_{\mu\nu}+h_{\mu\nu}$. This calculation is given in the appendix.
@@
The ten dimensional gauge transformation of the Fierz–Pauli action are given by
$H_{MN}\rightarrow H_{MN}+\bar{\nabla}_{M}\xi_{N}+\bar{\nabla}_{N}\xi_{M}.$ (50)
These can be decomposed into gauge transformations on the various fields, where the Weyl transformation also transforms the gauge parameter as $\xi_{\mu}\rightarrow\zeta^{-3}\xi_{\mu}$. Along with the $\xi_{i}\rightarrow\zeta\xi_{i}$, this gives the 4d gauge transformations as
$\begin{split}h_{\mu\nu}&\rightarrow h_{\mu\nu}+\nabla_{\mu}\xi_{\nu}+\nabla_{\nu}\xi_{\mu},\\ A_{\mu i}&\rightarrow A_{\mu i}+\partial_{\mu}\xi_{i},\\ \varphi&\rightarrow\varphi-\tfrac{4}{\sqrt{3}}\zeta^{-4}\xi_{0}.\end{split}$ (51)
The $\xi_{i}$ components result in gauge transformations for the gauge field and leave the $F_{i}^{2}$ term invariant. The gauge parameter $\xi_{\mu}$ corresponds to the active coordinate transformations $x^{\mu}\rightarrow x^{\mu}+\xi^{\mu}$, which act on the metric in the familiar way. They also act on the dilaton as expected in a kinating background. Under the active coordinate transformation the kination field can be related as $\Phi(x^{\mu})=\Phi^{\prime}(x^{\mu}-\xi^{\mu})$. If the dilaton is a small perturbation around a background kinating field $\Phi=\bar{\Phi}+\varphi$ then
$\bar{\Phi}(\eta)+\varphi(x^{\mu})=\bar{\Phi}(\eta-\xi^{0})+\tilde{\varphi}(x^{\mu}-\xi^{\mu}).$ (52)
Taylor expanding the background $\bar{\Phi}(\eta-\xi^{0})$, implies that the dilaton transformations as
$\varphi(x^{\mu})\rightarrow\varphi^{\prime}(x^{\mu})=\varphi(x^{\mu})-\zeta^{-2}\bar{\Phi}^{\prime}\xi_{0}.$ (53)
Since $\bar{\Phi}^{\prime}(\eta)=\tfrac{4}{\sqrt{3}}\zeta^{-2}$, this yields the same transformation as acquired from the compactification of the Kasner solution. In the appendix, we illustrate the explicit gauge invariance of 4d action for one of these transformations.